# Estimation of Optimal Causal Bounds via Covariate-Assisted Optimal Transport

### v1

@@ Appendix D Proofs of Section 4
Lemma D.7 (Linear functional gap).
@@
Under Assp. 2.3 and 4.6, we have
$\mathbb{E}[\mathds{W}_{1}(\hat{\mu}_{Z(0),n},\tilde{\mu}_{Z(0),n})]\leq\frac{\sqrt{d_{Z}}}{2\delta^{3}}n^{-\frac{1}{2}}+\frac{\sqrt{d_{Z}}C_{w}}{2\delta}n^{-r}.$
Lemma D.8 (Empirical Wasserstein convergence rate for reweighted empirical distribution).
Under the same condition of Lemma D.7, we have
$\mathbb{E}[\mathds{W}_{1}(\hat{\mu}_{Z(0),n},\mu_{Z(0),n})]\leq\delta^{-2}CR_{d_{Z}}(n)+\frac{\sqrt{d_{Z}}}{2\delta^{3}}n^{-\frac{1}{2}}+\frac{\sqrt{d_{Z}}C_{w}}{2\delta}n^{-r},$
@@
Lemma D.9 (Discretization gap).
Under Assp. 4.4 and 4.5 with
@@
where $\hat{w}_{i}$ is defined as in Lemma D.7, we have
@@
Lemma D.10 (Gap between $\hat{\bm{\mu}}$ and $\mu$).
Under Assp. 3.1 and 3.4 with $\hat{\bm{\mu}}_{Z(0),n}$ defined as in Lemma D.9, we have
@@
Proof of Lemma D.10.
The proof follows exactly the same as the proof of Lemma D.3, which holds as long as $\hat{\bm{\mu}}_{Z(0),n}$ is a probability measure supported on the cell centers $\{\varphi_{r}^{n}(G):G\in\Phi_{r}^{n}\}$. ∎
We also introduce the following theorem adapted from [backhoff2022estimating, Lemma 3.3].
Lemma D.11 (Conditional distribution of $Y$).
Under Assp. 4.3, conditionally on {|$\{|\{i:Z_{i}(0)\in G\}|,G\in\Phi_{r}^{n}\}$, for each $G$, the sample $(Y_{i}(0),i:Z_{i}(0)\in G)$ i.i.d. follows $\mu(\cdot|Z(0)\in G,W=0)$.
Proof of Theorem 4.3.
@@
The proof strategy of Theorem 4.1 is similar to that of Theorem 4.2.
@@
In the following, we adopt the same notation as in the proof of Theorem 4.2, omitting the dependence of $n,m$ in subscripts and superscripts for notational simplicity.
@@
$|\hat{V}_{m,n}-V_{\textup{c}}|\leq$ $\underbrace{L_{h}\left(\int\mathds{W}_{1}(\hat{\bm{\mu}}^{z_{0}}_{Y(0)},\mu^{z_{0}}_{Y(0)}){\rm d}\hat{\bm{\mu}}_{Z(0)}(z_{0})+\int\mathds{W}_{1}(\mu^{z_{1}}_{Y(1)},\hat{\bm{\mu}}^{z_{1}}_{Y(1)}){\rm d}\hat{\bm{\mu}}_{Z(1)}(z_{1})\right)}_{(\textup{Term A})}$
@@
For (Term C), similar to the derivation in Remark D.2, we have
@@
where the first inequality is the triangle inequality (Lemma A.1) and the second is due to Lemma D.9.
Under Assp. 2.3 and  4.6, by Lemma D.8, we get
$\mathbb{E}\left[\mathds{W}_{1}(\hat{\mu}_{Z(0)},\mu_{Z})\right]\leq\delta^{-2}CR_{d_{Z}}(n)+\sqrt{d_{Z}}C_{w}n^{-r}.$
This is due to $\frac{{\rm d}\mu}{{\rm d}\mu_{|W=0}}(z)=(1-e(z))^{-1}\mathbb{P}(W=0)\leq\delta^{-1}$ (Assp. 4.6). Combine these, we get
[…]
@@
By Lemma C.2 and Assp. 3.4, the mapping $z\mapsto\mathds{W}_{h}(\mu^{z}_{Y(0)},\mu^{z}_{Y(1)})$ is $(2L_{h}L_{Z})$-Lipschitz continuous. Then, by Lemma D.9,
@@
By Lemma D.7, and the fact that the mapping $z\mapsto\mathds{W}_{h}(\mu^{z}_{Y(0)},\mu^{z}_{Y(1)})$ is $(2L_{h}L_{Z})$-Lipschitz continuous, we get
(Term $(\text{Term E})\leq 2L_{h}L_{Z}(\frac{\sqrt{d_{Z}}}{2\delta^{3}}n^{-\frac{1}{2}}+\frac{\sqrt{d_{Z}}C_{w}}{2\delta}n^{-r}).$
And by Assp. 2.2, Lemma D.1 and Markov’s inequality, we get
(Term $(\text{Term F})\leq\delta^{-1}\left(16(L_{h}L_{Z})^{2}\int\|z\|_{2}^{2}{\rm d}\mu_{Z}(z)+C_{\mu,h}\right)^{\frac{1}{2}}n^{-\frac{1}{2}}\leq C^{\prime}_{2}n^{-r},$
@@
Let $\nu=\mu_{Z|W=0}\otimes\mu_{Y(0)}^{Z}$, thus by Assp. 2.2, $\nu^{z}_{Y(0)}=\mu^{z}_{Y(0)}$. Then, by Lemma D.3,
@@
Note that $(\hat{\bm{\mu}}_{Z(0)}(G),G\in\Phi_{r}^{n})$ only depends on $\hat{e}$ and $\mathcal{G}=\{|\{i:Z_{i}(0)\in G\}|,G\in\Phi_{r}^{n}\}$, then apply Lemma D.11, we get,
@@
$=$ $\mathbb{E}\left[\sum_{G\in\Phi_{r}^{n}}\hat{\bm{\mu}}_{Z(0)}(G)\mathbb{E}\left[\mathds{W}_{1}\left(\frac{1}{|\{i:Z_{i}(0)\in G\}|}\sum_{i:Z_{i}(0)\in G}\delta_{Y_{i}(0)},\nu_{Y(0)}^{G}\right)|\mathcal{G},\hat{e}\right]\right]$
@@
where the first inequality is also due to Lemma D.11 with independence between $\hat{e}$ and the sample $Z,Y$, the second is due to Assp. 4.5, and the last inequality is due to $x\mapsto xR_{d_{Y}}(x)$ is concave.
@@
$R_{d_{Y}}(\cdot)$ is defined in (10), and $C^{\prime}_{4}$ is a constant that depends on $d_{Y},M$ (Assp. 3.2). Therefore, we get
@@
Let $r=\frac{1}{d_{Z}+2\vee d_{Y}}$, the remaining steps follow the same as the proof of Theorem 4.2. ∎
Appendix E Proofs of Supporting Lemmas
E.1 Proof of Lemma A.1
Proof of Lemma A.1.
Let $\pi$ ($\tilde{\pi}$, resp.) be an optimal coupling associated with $\mathds{W}_{1}(\mu,\nu)$ ($\mathds{W}_{1}(\nu,\mu^{\prime})$, resp.). By the gluing lemma in Chapter 1 of [villani2009optimal], there is a coupling $\gamma=\gamma(x,y,x^{\prime})$ such that
@@ Appendix E Proofs of Supporting Lemmas
E.2 Proof of Lemma A.2
Proof of Lemma A.2.
@@
$\mathds{W}_{1}(\lambda\mu+(1-\lambda)\nu,\mu^{\prime})$ $\leq\int\left\|{x-x^{\prime}}\right\|_{{2}}{\rm d}(\lambda\pi+(1-\lambda)\tilde{\pi})$
@@
E.3 Proof of Lemma C.1
Proof of Lemma C.1.
Since $\Pi_{\textup{c}}\subseteq\Pi(\mu_{Y(0)},\mu_{Y(1)})$ and $\Pi(\mu_{Y(0)},\mu_{Y(1)})$ is tight ([villani2009optimal, Lemma 4.4]), then $\Pi_{\textup{c}}$ is also tight. It remains to show that $\Pi_{\textup{c}}$ is closed with respect to the weak topology. Suppose that $\pi_{k}\in\Pi_{\textup{c}}$ for $k\geq 1$ and $\lim_{k}\pi_{k}=\pi_{\infty}$, then for any bounded continuous function $\ell:\mathcal{Y}\times\mathcal{Z}\rightarrow\mathbb{R}$, $\mathbb{E}_{\pi_{\infty}}[\ell(Y(0),Z)]=\lim_{k}\mathbb{E}_{\pi_{k}}[\ell(Y(0),Z)]=\int\ell{\rm d}\mu_{Y(0),Z}$. Similarly, $\mathbb{E}_{\pi_{\infty}}[\ell(Y(1),Z)]=\int\ell{\rm d}\mu_{Y(1),Z}$. Therefore, $\pi_{\infty}\in\Pi_{\textup{c}}$, thus $\Pi_{\textup{c}}$ is closed. ∎
E.4 Proof of Lemma C.2
Proof of Lemma C.2.
@@
$\left|\mathds{W}_{h}(\mu,\nu)-\mathds{W}_{h}(\mu,\nu^{\prime})\right|\leq L_{h}\mathds{W}_{1}(\nu,\nu^{\prime}).$
@@
$\left|\mathds{W}_{h}(\mu,\nu)-\mathds{W}_{h}(\mu^{\prime},\nu^{\prime})\right|\leq\left|\mathds{W}_{h}(\mu,\nu)-\mathds{W}_{h}(\mu,\nu^{\prime})\right|+\left|\mathds{W}_{h}(\mu,\nu^{\prime})-\mathds{W}_{h}(\mu^{\prime},\nu^{\prime})\right|.$ (16)
@@
Let $\pi_{1}\in\mathcal{P}(\mathcal{Y}^{2})$ be the optimal coupling associated with $\mathds{W}_{h}(\mu,\nu)$, and let $\pi_{2}\in\mathcal{P}(\mathcal{Y}^{2})$ be the optimal coupling associated with $\mathds{W}_{1}\left(\nu,\nu^{\prime}\right)$. By the gluing lemma in Chapter 1 of [villani2009optimal], there is a coupling $\pi_{3}\in\mathcal{P}(\mathcal{Y}^{3})$, where we denote $\mathcal{Y}^{3}=\{(y^{(1)},y^{(2)},y^{(3)}),y^{(k)}\in\mathcal{Y}\}$, such that
@@
Therefore, by Assp. 3.3, we have
@@
$\leq$ $\mathbb{E}_{\pi_{3}}\left[h(y^{(1)},y^{(2)})+L_{h}\|y^{(2)}-y^{(3)}\|_{2}\right]$
$=$ $\mathds{W}_{h}(\mu,\nu)+L_{h}\mathds{W}_{1}\left(\nu,\nu^{\prime}\right).$
where the first inequality is due to $\pi_{3}(y^{(1)},y^{(3)})\in\Pi(z,z_{1})$; the second inequality is due to the Lipschitz continuity of $h$ by Assp. 3.3; the equation is due to the construction of $\pi_{3}$ by gluing $\pi_{1},\pi_{2}$.
@@
$\left|\mathds{W}_{h}(\mu,\nu)-\mathds{W}_{h}(\mu,\nu^{\prime})\right|\leq L_{h}\mathds{W}_{1}\left(\nu,\nu^{\prime}\right).$
@@
E.5 Proof of Lemma D.1
Proof of Lemma D.1.
@@
For (Term A), by Lemma C.2 and Assp. 3.4, we have
@@
E.6 Proof of Lemma D.2
Proof of Lemma D.2.
@@
E.7 Proof of Lemma D.3
Lemma E.1.
@@
Then, by Lemma E.1, we have
@@
By the convexity of $\mu\mapsto\mathds{W}_{1}(\mu,\mu_{Y(0)}^{\zeta})$ (Lemma A.2) and Assp. 3.4, for $\zeta\in\varphi_{r}^{n}([0,1]^{d_{Z}})$, we have
$\mathds{W}_{1}(\mu^{G}_{Y(0)},\mu^{\zeta}_{Y(0)})=$ $\mathds{W}_{1}\left(\int_{G}\mu^{\zeta^{\prime}}_{Y(0)}{\rm d}\mu_{Z}(\zeta^{\prime}),\mu^{\zeta}_{Y(0)}\right)$
@@
Back to (17), by Lemma E.1 that $\hat{\mu}_{Z(0)}(G)=\hat{\bm{\mu}}_{Z(0)}(G)$, we obtain
@@
E.8 Proof of Lemma D.6
Proof of Lemma D.6.

### v2

@@ F.4 Proof of Theorem 6
Lemma 12 (Linear functional gap).
@@
Under Assp. 3 and 13, we have
$\mathbb{E}[\mathds{W}_{1}(\hat{\mu}_{Z(0),n},\tilde{\mu}_{Z(0),n})]\leq\frac{\sqrt{d_{Z}}}{2\delta^{3}}n^{-\frac{1}{2}}+\frac{\sqrt{d_{Z}}C_{\textup{w}}}{2\delta}n^{-r}.$
Lemma 13 (Empirical Wasserstein convergence rate for reweighted empirical distribution).
Under the same condition of Lemma 12, we have
$\mathbb{E}[\mathds{W}_{1}(\hat{\mu}_{Z(0),n},\mu_{Z(0),n})]\leq\delta^{-2}CR_{d_{Z}}(n)+\frac{\sqrt{d_{Z}}}{2\delta^{3}}n^{-\frac{1}{2}}+\frac{\sqrt{d_{Z}}C_{\textup{w}}}{2\delta}n^{-r},$
@@
Lemma 14 (Discretization gap).
Under Assp. 11 and 12 with
@@
where $\hat{w}_{i}$ is defined as in Lemma 12, we have
@@
Lemma 15 (Gap between $\hat{\bm{\mu}}$ and $\mu$).
Under Assp. 4 and 8 with $\hat{\bm{\mu}}_{Z(0),n}$ defined as in Lemma 14, we have
@@
Proof of Lemma 15.
The proof follows exactly the same as the proof of Lemma 8, which holds as long as $\hat{\bm{\mu}}_{Z(0),n}$ is a probability measure supported on the cell centers $\{\varphi_{r}^{n}(G):G\in\Phi_{r}^{n}\}$. ∎
We also introduce the following theorem adapted from [3, Lemma 3.3].
Lemma 16 (Conditional distribution of $Y$).
Under Assp. 10, conditionally on $\{|\{i:Z_{i}(0)\in G\}|,G\in\Phi_{r}^{n}\}$, for each $G$, the sample $(Y_{i}(0),i:Z_{i}(0)\in G)$ i.i.d. follows $\mu(\cdot|Z(0)\in G,W=0)$.
Proof of Theorem 6.
@@
The proof strategy of Theorem 3 is similar to that of Theorem 4.
@@
In the following, we adopt the same notation as in the proof of Theorem 4, omitting the dependence of $n,m$ in subscripts and superscripts for notational simplicity.
@@
$|\hat{V}_{n,m}-V_{\textup{c}}|\leq$ $\underbrace{L_{h}\left(\int\mathds{W}_{1}(\hat{\bm{\mu}}^{z_{0}}_{Y(0)},\mu^{z_{0}}_{Y(0)}){\rm d}\hat{\bm{\mu}}_{Z(0)}(z_{0})+\int\mathds{W}_{1}(\mu^{z_{1}}_{Y(1)},\hat{\bm{\mu}}^{z_{1}}_{Y(1)}){\rm d}\hat{\bm{\mu}}_{Z(1)}(z_{1})\right)}_{(\textup{Term A})}$
@@
For (Term C), similar to the derivation in Remark 4, we have
@@
where the first inequality is the triangle inequality (Lemma 1) and the second is due to Lemma 14.
Under Assp. 3 and  13, by Lemma 13, we get
$\mathbb{E}\left[\mathds{W}_{1}(\hat{\mu}_{Z(0)},\mu_{Z})\right]\leq\delta^{-2}CR_{d_{Z}}(n)+\sqrt{d_{Z}}C_{\textup{w}}n^{-r}.$
This is due to $\frac{{\rm d}\mu}{{\rm d}\mu_{|W=0}}(z)=(1-e(z))^{-1}\mathbb{P}(W=0)\leq\delta^{-1}$ (Assp. 13). Combine these, we get
$\text{(Term C)}\leq C^{\prime}_{1}(R_{d_{Z}}(n\wedge m)+(n\wedge m)^{-r}).$
@@
[…]
@@
[…]
And by Assp. 2, Lemma 6 and Markov’s inequality, we get
[…]
@@
Let $\nu=\mu_{Z|W=0}\otimes\mu_{Y(0)}^{Z}$, thus by Assp. 2, $\nu^{z}_{Y(0)}=\mu^{z}_{Y(0)}$. Then, by Lemma 8,
@@
Note that $(\hat{\bm{\mu}}_{Z(0)}(G),G\in\Phi_{r}^{n})$ only depends on $\hat{e}$ and $\mathcal{G}=\{|\{i:Z_{i}(0)\in G\}|,G\in\Phi_{r}^{n}\}$, then apply Lemma 16, we get,
@@
[…]
@@
where the first inequality is also due to Lemma 16 with independence between $\hat{e}$ and the sample $Z,Y$, the second is due to Assp. 12, and the last inequality is due to $x\mapsto xR_{d_{Y}}(x)$ is concave.
@@
$R_{d_{Y}}(\cdot)$ is defined in (8), and $C^{\prime}_{4}$ is a constant that depends on $d_{Y}$ and $\max_{y\in\mathcal{Y}}\left\|{y}\right\|_{{2}}$. Therefore, we get
@@
Let $r=\frac{1}{d_{Z}+2\vee d_{Y}}$, the remaining steps follow the same as the proof of Theorem 4. ∎
Appendix G Proof of Section 5
Lemma 17 ($V_{\textup{c}}$ for Gaussian noise model, [35, Lemma 6.1]).
  1. (i)
  2. (ii)
Appendix H Proofs of Supporting Lemmas
H.1 Proof of Lemma 1
Proof of Lemma 1.
Let $\pi$ ($\tilde{\pi}$, resp.) be an optimal coupling associated with $\mathds{W}_{1}(\mu,\nu)$ ($\mathds{W}_{1}(\nu,\mu^{\prime})$, resp.). By the gluing lemma in Chapter 1 of [49], there is a coupling $\gamma=\gamma(x,y,x^{\prime})$ such that
@@ H.1 Proof of Lemma 1
H.2 Proof of Lemma 2
Proof of Lemma 2.
@@ H.2 Proof of Lemma 2
[…]
@@
H.3 Proof of Lemma 3
Proof of Lemma 3.
Since $\Pi_{\textup{c}}\subseteq\Pi(\mu_{Y(0)},\mu_{Y(1)})$ and $\Pi(\mu_{Y(0)},\mu_{Y(1)})$ is tight ([49, Lemma 4.4]), then $\Pi_{\textup{c}}$ is also tight. It remains to show that $\Pi_{\textup{c}}$ is closed with respect to the weak topology. Suppose that $\pi_{k}\in\Pi_{\textup{c}}$ for $k\geq 1$ and $\lim_{k}\pi_{k}=\pi_{\infty}$, then for any bounded continuous function $\ell:\mathcal{Y}\times\mathcal{Z}\rightarrow\mathbb{R}$, $\mathbb{E}_{\pi_{\infty}}[\ell(Y(0),Z)]=\lim_{k}\mathbb{E}_{\pi_{k}}[\ell(Y(0),Z)]=\int\ell{\rm d}\mu_{Y(0),Z}$. Similarly, $\mathbb{E}_{\pi_{\infty}}[\ell(Y(1),Z)]=\int\ell{\rm d}\mu_{Y(1),Z}$. Therefore, $\pi_{\infty}\in\Pi_{\textup{c}}$, thus $\Pi_{\textup{c}}$ is closed. ∎
H.4 Proof of Lemma 4
Proof of Lemma 4.
On the compact space $\mathcal{Y}$, the weak topology can be induced by the metric $\mathds{W}_{1}$, thus $g_{w}(z)=\mu_{Y(w)}^{z}\ w=0,1$ are continuous with respect to $\mathds{W}_{1}$. Since $\mathcal{Z}$ is compact, then $g_{w}(z)\ w=0,1$ are uniformly continuous with respect to $\mathds{W}_{1}$. As a result, we get that, for any $\delta>0$, there is a constant $C(\delta)$ such that $\mathds{W}_{1}\left(\mu_{Y(k)}^{z_{0}},\mu_{Y(k)}^{z_{1}}\right)\leq\delta+C(\delta)\|z_{0}-z_{1}\|_{2}$. ∎
H.5 Proof of Lemma 5
Proof of Lemma 5.
@@ H.5 Proof of Lemma 5
$\left|\mathds{W}_{h}(\mu,\nu)-\mathds{W}_{h}(\mu,\nu^{\prime})\right|\leq L_{h}\mathds{W}_{1}(\nu,\nu^{\prime})+\epsilon.$
@@
$\left|\mathds{W}_{h}(\mu,\nu)-\mathds{W}_{h}(\mu^{\prime},\nu^{\prime})\right|\leq\left|\mathds{W}_{h}(\mu,\nu)-\mathds{W}_{h}(\mu,\nu^{\prime})\right|+\left|\mathds{W}_{h}(\mu,\nu^{\prime})-\mathds{W}_{h}(\mu^{\prime},\nu^{\prime})\right|+2\epsilon.$ (17)
@@
Let $\pi_{1}\in\mathcal{P}(\mathcal{Y}^{2})$ be the optimal coupling associated with $\mathds{W}_{h}(\mu,\nu)$, and let $\pi_{2}\in\mathcal{P}(\mathcal{Y}^{2})$ be the optimal coupling associated with $\mathds{W}_{1}\left(\nu,\nu^{\prime}\right)$. By the gluing lemma in Chapter 1 of [49], there is a coupling $\pi_{3}\in\mathcal{P}(\mathcal{Y}^{3})$, where we denote $\mathcal{Y}^{3}=\{(y^{(1)},y^{(2)},y^{(3)}),y^{(k)}\in\mathcal{Y}\}$, such that
@@
Therefore, by Assp. 7, we have
@@
$\leq$ $\mathbb{E}_{\pi_{3}}\left[h(y^{(1)},y^{(2)})+L_{h}\|y^{(2)}-y^{(3)}\|_{2}+\epsilon\right]$
$=$ $\mathds{W}_{h}(\mu,\nu)+L_{h}\mathds{W}_{1}\left(\nu,\nu^{\prime}\right)+\epsilon.$
where the first inequality is due to $\pi_{3}(y^{(1)},y^{(3)})\in\Pi(z,z_{1})$; the second inequality is due to the smoothness assumption of $h$ (10); the equation is due to the construction of $\pi_{3}$ by gluing $\pi_{1},\pi_{2}$.
@@
$\left|\mathds{W}_{h}(\mu,\nu)-\mathds{W}_{h}(\mu,\nu^{\prime})\right|\leq L_{h}\mathds{W}_{1}\left(\nu,\nu^{\prime}\right)+\epsilon.$
@@
H.6 Proof of Lemma 6
Proof of Lemma 6.
@@ H.6 Proof of Lemma 6
For (Term A), by Lemma 5 and Assp. 8, we have
@@
H.7 Proof of Lemma 7
Proof of Lemma 7.
@@ H.7 Proof of Lemma 7
H.8 Proof of Lemma 8
Lemma 18.
@@ H.8 Proof of Lemma 8
Then, by Lemma 18, we have
$\mathds{W}_{1}(\hat{\bm{\mu}}_{Y(0)}^{G},\mu_{Y(0)}^{\zeta})\leq$ $\mathds{W}_{1}(\hat{\bm{\mu}}_{Y(0)}^{G},\mu_{Y(0)}^{G})+\mathds{W}_{1}(\mu_{Y(0)}^{G},\mu_{Y(0)}^{\zeta})$ $(\textup{by the triangle ineq.})$
@@
By the convexity of $\mu\mapsto\mathds{W}_{1}(\mu,\mu_{Y(0)}^{\zeta})$ (Lemma 2) and Assp. 8, for $\zeta\in\varphi_{r}^{n}([0,1]^{d_{Z}})$, we have
@@
Back to (18), by Lemma 18 that $\hat{\mu}_{Z(0)}(G)=\hat{\bm{\mu}}_{Z(0)}(G)$, we obtain
@@
H.9 Proof of Lemma 11
Proof of Lemma 11.